# COMPASS: Robust Feature Conformal Prediction for Medical Segmentation Metrics

### v1

@@ Appendix E Experimental Details
We use the standard U-Net, SegResNet, and SwinUNETR architecture available in the MONAI framework [25]. For the U-Net, we used an encoder with channel sizes of (32, 64, 128, 256) and a corresponding decoder, using two residual units per block, batch normalization, and a dropout rate of 0.1. For SegResNet, we used a residual encoder-decoder network that started with 32 initial filters and included varying numbers of blocks in its down-sampling path (1, 2, 2, 4), batch normalization, and a dropout probability of 0.1. For SwinUNETR, we used a transformer-based model configured for an image size of (128, 128) pixels, with a feature size of 48 and multi-head attention mechanisms across four depth levels.
@@
For each architecture, we trained both a standard model and an End-to-End Conformalized Quantile Regression (E2E-CQR) variant [24], with all models being trained using an AdamW optimizer with a learning rate of 1e-4 and a batch size of 32. The standard models were configured for binary segmentation tasks, taking a 1 (for grayscale) or 3-channel (for RGB image) as input and producing a single-channel output mask. Training was optimized using a Dice score loss function to maximize the overlap between the predicted and ground-truth segmentations. Following the original E2E-CQR model [24], we modify the number of output channels to 3 to produce three distinct segmentation masks, corresponding to the lower quantile, the median prediction, and the upper quantile, which together form the prediction interval. These models were trained using the Tversky loss [33]. Thus, it allows the model to directly learn the uncertainty bounds end-to-end.
@@
We follow the standard frameworks for SCP [8, 29], CQR [30], and Local CP [31, 32, 29]. For CQR, we use two separate Gradient Boosting Regressors to learn the lower and upper quantiles of the prediction and adjust the resulting We use a learning rate of 0.1, 50 estimators, a maximum depth of 3, minimum leaf samples of 1, and a minimum samples to split of 9. For Local CP, we use two separate Random Forest Regressors to learn the mean and mean absolute difference. We use 1000 estimators and a minimum leaf sample of 100. For E2E-CQR, we follow the same calibration procedure as CQR. However, in this case, the quantiles are learned by the model.
@@
For all experiments, we run both the symmetric and asymmetric versions of COMPASS with 1 component on the logits and each layer in the segmentation head. We report the layer with the minimum mean interval length and its corresponding coverage. We perform 100 random calibration-test splits. For weighted CP, we use LightGBM with the default settings [37]. We also perform 100 calibration-test splits. However, in this case, the samples were shuffled to maintain the target proportion of classes. We show the covariate shift in Figure 5.
Appendix F $\beta$-$\delta A$ Trajectories
@@
Figure 6: Figure perturbation trajectories for 4 datasets for U-Net
Appendix G Additional Figures
@@
Figure 7: Segmentation volume increases with $\beta$ for H&E. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 8: Segmentation volume increases with $\beta$ for Skin Lesion. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 9: Segmentation volume increases with $\beta$ for Thyroid Nodule. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 10: Segmentation volume increases with $\beta$ for PolyP. Each row is a different sample. $\delta A=0\%$ is the original prediction.
@@
Figure 11: Segmentation volume increases with $\beta$ for H&E. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 12: Segmentation volume increases with $\beta$ for Skin Lesion. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 13: Segmentation volume increases with $\beta$ for Thyroid Nodule. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 14: Segmentation volume increases with $\beta$ for PolyP. Each row is a different sample. $\delta A=0\%$ is the original prediction.

### v2

@@ Appendix E Experimental Details
We use the standard U-Net, SegResNet, and SwinUNETR architecture available in the MONAI framework [Cardoso et al., 2022]. For the U-Net, we used an encoder with channel sizes of (32, 64, 128, 256) and a corresponding decoder, using two residual units per block, batch normalization, and a dropout rate of 0.1. For SegResNet, we used a residual encoder-decoder network that started with 32 initial filters and included varying numbers of blocks in its down-sampling path (1, 2, 2, 4), batch normalization, and a dropout probability of 0.1. For SwinUNETR, we used a transformer-based model configured for an image size of (128, 128) pixels, with a feature size of 48 and multi-head attention mechanisms across four depth levels.
@@
For each architecture, we trained both a standard model and an End-to-End Conformalized Quantile Regression (E2E-CQR) variant [Lambert et al., 2024], with all models being trained using an AdamW optimizer with a learning rate of 1e-4 and a batch size of 32. The standard models were configured for binary segmentation tasks, taking a 1 (for grayscale) or 3-channel (for RGB image) as input and producing a single-channel output mask. Training was optimized using a Dice score loss function to maximize the overlap between the predicted and ground-truth segmentations. Following the original E2E-CQR model [Lambert et al., 2024], we modify the number of output channels to 3 to produce three distinct segmentation masks, corresponding to the lower quantile, the median prediction, and the upper quantile, which together form the prediction interval. These models were trained using the Tversky loss [Salehi et al., 2017]. Thus, it allows the model to directly learn the uncertainty bounds end-to-end.
@@
We follow the standard frameworks for SCP [Papadopoulos et al., 2002, Lei et al., 2018], CQR [Romano et al., 2019], and Local CP [Papadopoulos et al., 2008, 2011, Lei et al., 2018]. For CQR, we use two separate Gradient Boosting Regressors to learn the lower and upper quantiles of the prediction and adjust the resulting We use a learning rate of 0.1, 50 estimators, a maximum depth of 3, minimum leaf samples of 1, and a minimum samples to split of 9. For Local CP, we use two separate Random Forest Regressors to learn the mean and mean absolute difference. We use 1000 estimators and a minimum leaf sample of 100. For E2E-CQR, we follow the same calibration procedure as CQR. However, in this case, the quantiles are learned by the model.
@@
For all experiments, we run both the symmetric and asymmetric versions of COMPASS with 1 component on the logits and each layer in the segmentation head. We report the layer with the minimum mean interval length and its corresponding coverage. We perform 100 random calibration-test splits. For weighted CP, we use LightGBM with the default settings [Ke et al., 2017]. We also perform 100 calibration-test splits. However, in this case, the samples were shuffled to maintain the target proportion of classes. We show the covariate shift in Figure 6.
Appendix F Practical Recommendation Additional Results
Figure 7: COMPASS results in stable calibration. We use the results from $\alpha$=0.1, binned the ground truth values into 5 non-overlapping bins, and computed the mean interval length and mean coverage (and their 95% confidence intervals). We find that COMPASS methods tend to perform better for shorter interval lengths (except for H&E), achieving approximately the same coverage. The count of the bins tends to be a better indicator of calibration stability (See Figure 8). Please note that our guarantees are only marginal; therefore, coverage may be lower for values that are less frequently represented during calibration.
Figure 8: Ground Truth Distribution. We show histogram plots of the distribution of ground truth segmentation area values in Pixels2 with a KDE overlaid.
Figure 9: One principal component is generally sufficient for COMPASS-J. For 4 datasets, we run COMPASS-J with 1 to 5 components for 100 iterations. We find that there is no additional benefits of using more than 1 principal component, as the first principal component already explained most of the variance (Figure 5).
Figure 10: Explained variance of the Jacobians is significantly lower when computed on raw flattened features. For our 4 datasets, we plot the first principal component’s explained variance against the feature layer used for COMPASS-J. We find all layers result in non-monotonicity, indicating the translation-variant nature of the flattened space.
Table 8: Jacobian computation for COMPASS-J is fast. We present the average computational time to compute the Jacobians of feature size 64x64x64 across four datasets below using a single NVIDIA A100 GPU. We compute the Jacobians with respect to the metric, which is extremely fast with autograd. This is because the downstream metric is one-dimensional. Furthermore, to expedite the training and calibration step, we precompute and save the Jacobians summed on the spatial dimension.
H&E Skin Lesion Nodule PolyP
Compute time for full dataset
(Mean±STD in seconds)
73.6±0.6 305.8±1.1 93.3±0.9 73.6±0.6
Figure 11: COMPASS methods do not require substantial computing times. The calibration step requires multiple forward passes. However, we find that combining forward passes and binary search enables fast calibration. We show the average computational time required for calibration (for 100 repeats) across 4 datasets using a single NVIDIA A100 GPU. We compute times for the symmetric (without A-) and asymmetric (A-), deep (-Deep) and (-Shallow) layers, and COMPASS-L versus COMPASS-J. The full calibration runtimes are on the order of seconds for the full calibration dataset. The number of samples used for calibration was 223 (H&E), 1000 (Skin Lesion), 349 (Thyroid), and 200 (PolyP).
@@ Appendix F Practical Recommendation Additional Results
Appendix G Segmentation Overlays for COMPASS-J and COMPASS-L
@@
Figure 13: Segmentation area increases with $\beta$ for H&E. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 14: Segmentation area increases with $\beta$ for Skin Lesion. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 15: Segmentation area increases with $\beta$ for Thyroid Nodule. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 16: Segmentation area increases with $\beta$ for PolyP. Each row is a different sample. $\delta A=0\%$ is the original prediction.
@@
Figure 17: Segmentation area increases with $\beta$ for H&E. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 18: Segmentation area increases with $\beta$ for Skin Lesion. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 19: Segmentation area increases with $\beta$ for Thyroid Nodule. Each row is a different sample. $\delta A=0\%$ is the original prediction.
Figure 20: Segmentation area increases with $\beta$ for PolyP. Each row is a different sample. $\delta A=0\%$ is the original prediction.